OpenStudy (anonymous):

use an algebraic equation to find the measures of the two angles described below. Begin by letting x represent the degree measure of the angle's supplement. The measure of the angle is nine times greater than its supplement What is the measure of the supplement? What is the measure of the other angle?

OpenStudy (shamil98):

Supplementary angles add up to 180. so x + y = 180 9x = y

OpenStudy (shamil98):

Substitution, x + 9x = 180 10x = 180 Solve for x from there.

OpenStudy (anonymous):

18?

OpenStudy (shamil98):

mhm, then substitute x and solve for y 18 + y = 180

OpenStudy (anonymous):

162?

OpenStudy (shamil98):

yes, y = 162 and x = 18

OpenStudy (anonymous):

What is the measure of the supplement? What is the measure of the other angle? So what would i put for these questions

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

Begin by letting x represent the degree measure of the angle's supplement. so 9x = y The measure of the angle is nine times greater than its supplement the angle's supplement is y, the other angle is x.

OpenStudy (anonymous):

ohh okay!! Thank you so much!